Compound x is made of two parts A and one part B. The mass of compound X is 45, and A has twice the mass of B. Which two equations would you solve to find the individual masses of A and B?
1
Expert's answer
2021-10-05T01:31:03-0400

Compound x is made of two parts A and one part B.

Hence, the compound x is 2A+B

Let’s use the individual mass of A is "m_A" and individual mass of B is "m_B" .

The mass of compound X is 45

"2m_A + m_B=45 \\

m_A=2m_B \\

m_A -2m_B=0 \\

2m_A -4m_B =0 \\

2m_A -2m_A +m_B + 4m_B = 45 \\

m_B = \frac{45}{5} = 9 \\

m_A = 2 \times 9 = 18"